How many molecules are there in 1.00g of NH2


1
Expert's answer
2020-08-07T14:19:24-0400

M(NH2) = 16.02 g/mol

n=m/M

n(NH2) = 1 g / 16.02 (g/mol) = 0.0624 mol

One mole of NH2 molecules contains 6.02×1023 NH2 molecules.

Number of molecules = n×6.02×1023 = 0.0624×6.02×1023 = 0.375×1023 molecules.

Answer: 3.75×1022 molecules